OpenStudy (anonymous):

so I'm in the military which means that I been out of school for a while...just need a re-fresher with this problem (3/2 x3 1/2) divided by (6/5-1)

OpenStudy (anonymous):

okay is it \[ \large{\frac{\frac{3}{2} \times 3\frac{1}{2}}{\frac{6}{5}-1}}\]

OpenStudy (anonymous):

105/2

OpenStudy (anonymous):

so I know I was do PEMDAS right

OpenStudy (anonymous):

no...has nothing to do with order of operations

OpenStudy (anonymous):

solve the top first....3/2 x 7/2 is 21/4

OpenStudy (anonymous):

then solve the bottom is 1/5

OpenStudy (anonymous):

so you get 21/4 / 1/5....which means you flip and multiply which gives you 105/4

OpenStudy (anonymous):

thank you sooooooo much...just got it

OpenStudy (anonymous):

about to go show off to my fellow shipmates now....:0)